# Gnothi Seauton: Empowering Faithful Self-Interpretability in Black-Box Models

### v1

@@ 4.1 Efficiently Training the Shapley Value Explainer for Black-Box Models
To improve training, memory, and inference efficiency, we propose a side-tuning pipeline called AutoGnothi, as shown in Figure 3(b). Building on ideas PETL, we adapt Ladder Side-Tuning (LST) (Sung et al., 2022) by incorporating an additive side network. This side network separates the trainable parameters from the backbone model $f$ and adapts the model to a different task. It is a lightweight version of $f$, with weights and hidden state dimensions scaled by a factor of $1/r$, where $r$ is a reduction factor (e.g., $r=4$ or $8$). For instance, if the backbone $f$ has a $768$-dimensional hidden state, then with $r=8$, the side network has a hidden state dimension of $96$. By computing gradients solely for the side network, this design avoids a backward pass through the main backbone, improving training and memory efficiency. The formulation combines the frozen pre-trained backbone and the side-tuner with learnable parameters $\beta$ as:
@@ 5.3 Evaluating Explanation Quality
Qualitative Results. We also provide visualization results for ViTs on different datasets, as shown in Figure 7. It may be observed that RISE happened to fail to provide human-interpretable or intuitive results, which amongst all others AutoGnothi offers more accurate explanations with a clearer focus on the subject and diluted colours for irrelevant classes. Additional visualization results for ViT-base on more datasets are provided in Appendix D, shown in Figures 89101112, and 13.
@@ 6 Conclusion
This paper introduces AutoGnothi to bridge the gap between self-interpretable models and post-hoc explanation methods in Explainable AI. Inspired by parameter-efficient transfer-learning, AutoGnothi incorporates a lightweight side network that allows black-box models to generate faithful Shapley value explanations without affecting the original predictions. Notably, AutoGnothi outperforms directly finetuning the all the parameters in the model for explanation by a clear margin. This approach empowers black-box models with self-interpretability, which is superior to standard post-hoc explanations that require generating predictions and explanations in two separate, heavy inferences. Experiments on ViT and BERT demonstrate that AutoGnothi achieves superior efficiency on computation, storage and memory in both training and inference periods.
@@ Appendix B Proofs of Theorems
$=Q^{-1}\sum_{k=1}^{d-1}\frac{d-1}{d(d-k)}$
$=\frac{\sum_{k=1}^{d-1}\frac{d-1}{d(d-k)}}{\sum_{k=1}^{d-1}\frac{d-1}{k(d-k)}}$
@@
$=Q^{-1}\sum_{k=2}^{d-1}\frac{k-1}{d(d-k)}$
$=\frac{\sum_{k=2}^{d-1}\frac{k-1}{d(d-k)}}{\sum_{k=1}^{d-1}\frac{d-1}{k(d-k)}}.$
@@
$=\frac{\frac{1}{d}+\sum_{k=2}^{d-1}\frac{d-k}{d(d-k)}}{\sum_{k=1}^{d-1}\frac{d-1}{k(d-k)}}$
$=\frac{\frac{1}{d}+\frac{d-2}{d}}{\sum_{k=1}^{d-1}\frac{d-1}{k(d-k)}}$
$=\frac{d-1}{d}\cdot\frac{1}{\sum_{k=1}^{d-1}\frac{d-1}{k(d-k)}}$
$=\frac{1}{2\sum_{k=1}^{d-1}\frac{1}{k}}$
$=\frac{1}{2H_{d-1}}.$
@@ Appendix C Additional Results for AutoGnothi
Table 8: Performance metrics for ViT-base on MURA. MURA is a binary classification dataset, we calculate metrics for each of its two categories.
@@
Appendix D Additional Visualizations for AutoGnothi

### v2

# Gnothi Seauton: Empowering Faithful Self-Interpretability in Black-Box Transformers

@@ 4.1 Efficiently Training the Shapley Value Explainer for Black-Box Models
To improve training, memory, and inference efficiency, we propose a side-tuning pipeline called AutoGnothi, as shown in Figure 3(b). Building on ideas from PETL, we adapt Ladder Side-Tuning (LST) (Sung et al., 2022) by incorporating an additive side network. This side network separates the trainable parameters from the backbone model $f$ and adapts the model to a different task. It is a lightweight version of $f$, with weights and hidden state dimensions scaled by a factor of $1/r$, where $r$ is a reduction factor (e.g., $r=4$ or $8$). For instance, if the backbone $f$ has a $768$-dimensional hidden state, then with $r=8$, the side network has a hidden state dimension of $96$. By computing gradients solely for the side network, this design avoids a backward pass through the main backbone, improving training and memory efficiency. The formulation combines the frozen pre-trained backbone and the side-tuner with learnable parameters $\beta$ as:
@@ 5.3 Evaluating Explanation Quality
Qualitative Results. We also provide visualization results for ViTs on different datasets, as shown in Figure 7. It may be observed that RISE happened to fail to provide human-interpretable or intuitive results, which amongst all others AutoGnothi offers more accurate explanations with a clearer focus on the subject and diluted colours for irrelevant classes. Additional visualization results for ViT-base on more datasets are provided in Appendix E, shown in Figures 89101112, and 13.
@@ 6 Conclusion
This paper introduces AutoGnothi to bridge the gap between self-interpretable models and post-hoc explanation methods in Explainable AI. Inspired by parameter-efficient transfer-learning, AutoGnothi incorporates a lightweight side network that allows black-box models to generate faithful Shapley value explanations without affecting the original predictions. Notably, AutoGnothi outperforms directly fine-tuning the all the parameters in the model for explanation by a clear margin. This approach empowers black-box models with self-interpretability, which is superior to standard post-hoc explanations that require generating predictions and explanations in two separate, heavy inferences. Experiments on ViT and BERT demonstrate that AutoGnothi achieves superior efficiency on computation, storage and memory in both training and inference periods.
@@ Appendix B Proofs of Theorems
$=Q^{-1}\sum_{k=1}^{d-1}\frac{d-1}{d(d-k)}=\frac{\sum_{k=1}^{d-1}\frac{d-1}{d(d-k)}}{\sum_{k=1}^{d-1}\frac{d-1}{k(d-k)}}$
@@
$=Q^{-1}\sum_{k=2}^{d-1}\frac{k-1}{d(d-k)}=\frac{\sum_{k=2}^{d-1}\frac{k-1}{d(d-k)}}{\sum_{k=1}^{d-1}\frac{d-1}{k(d-k)}}.$
@@
$=\frac{\frac{1}{d}+\sum_{k=2}^{d-1}\frac{d-k}{d(d-k)}}{\sum_{k=1}^{d-1}\frac{d-1}{k(d-k)}}=\frac{\frac{1}{d}+\frac{d-2}{d}}{\sum_{k=1}^{d-1}\frac{d-1}{k(d-k)}}$
$=\frac{d-1}{d}\cdot\frac{1}{\sum_{k=1}^{d-1}\frac{d-1}{k(d-k)}}=\frac{1}{2\sum_{k=1}^{d-1}\frac{1}{k}}=\frac{1}{2H_{d-1}}.$
@@ Appendix C Additional Results for AutoGnothi
Table 8: Performance metrics for ViT-base on MURA. MURA is a binary classification dataset, we calculated metrics for each of its two categories.
@@
Appendix D Ablation Study on the Size of Side-Networks
In this section, we present an ablation study on the architecture and size of side-networks in AutoGnothi. Specifically, we investigated the effect of varying the reduction factor ($r=2,4,8,16,32$) on the side-networks of ViT-base model. To evaluate the impact, we compared the performance of AutoGnothi with Covert et al. (Covert et al., 2022) on the ImageNette dataset. The evaluation was performed using the insertion and deletion metrics, and the results are summarized in Table 10.
This study aimed to assess the balance between computational efficiency and explanation quality under different configurations of the side-network. The findings are as follows:
  • Too Large $r$: When the reduction factor $r$ was too large (resulting in a very small side-network), the network lacked sufficient capacity to learn the explanations effectively. As a result, the explanation quality degraded, with the side-network failing to capture enough information from the main branch.
  • Too Small $r$: Conversely, when $r$ was too small (resulting in a larger side-network), the network consumed excessive computational resources and interfered with the shared features of the predictor. This interference reduced the feature similarity between the prediction and explanation tasks, negatively impacting the faithfulness of the explainer.
  • Moderate $r$: A moderate reduction factor (e.g., $r=8$) provided the optimal trade-off between computational efficiency and explanation quality. This configuration allowed the side-network to effectively utilize the features from the predictor while maintaining resource efficiency.
These results highlight the importance of carefully selecting the reduction factor to optimize the performance of AutoGnothi across different transformer architectures. The identified trade-offs ensure that the side-network remains both effective and efficient, making it suitable for various applications.
Table 10: Explanation Quality Metrics of the ViT-base Explainer on ImageNette dataset.
Method Insertion (↑) Deletion (↓)
Covert et al. 0.9839$\pm$0.0375 0.8121$\pm$0.1768
AutoGnothi ($r$=2) 0.9894$\pm$0.0329 0.8687$\pm$0.1806
AutoGnothi ($r$=4) 0.9857$\pm$0.0251 0.7846$\pm$0.1957
AutoGnothi ($r$=8) 0.9874$\pm$0.0265 0.7954$\pm$0.2294
AutoGnothi ($r$=16) 0.9720$\pm$0.0288 0.6202$\pm$0.2023
AutoGnothi ($r$=32) 0.9520$\pm$0.0443 0.5547$\pm$0.1941
Appendix E Additional Visualizations for AutoGnothi